OpenStudy (anonymous):

Simplify the radical. the square root of 200 A. 2 and the square root of 10 B. 10 and the square root of 2 C. 100 and the square root of 2 D. 2 and the square root of 100

OpenStudy (anonymous):

\[\sqrt{200} = \sqrt{2*100} = 10*\sqrt{2}\]

OpenStudy (anonymous):

B

OpenStudy (anonymous):

U STUDY IN HIGH SCHOOL?